# A New Algorithm for Zero-Sum Linear-Quadratic Stochastic Differential Games in Infinite Horizons

### v1

@@ 2.2 Zero-Sum Linear-Quadratic Stochastic Differential Games in Infinite Horizons
2.3 Algorithm Design
In this paper, we aim to solve for the closed-loop saddle points numerically for the $({SDG})^{\,0}_{\infty}$ problem with given parameter values. Consider the sequences $\{P^{(k)}\}_{k\geq 0}$ and $\{Z^{(k)}\}_{k\geq 0}$ constructed according to the following procedure:
  1. 1.
    For $k=0$, set $P^{(0)}=0$, and let $Z^{(0)}$ be the unique stabilizing solution to the following algebraic Riccati equation with a definite quadratic term:
    $0=Z^{(0)}A_{(0)}+A_{(0)}^{\top}Z^{(0)}+\sum_{l=1}^{r}C_{l,(0)}^{\top}Z^{(0)}C_{l,(0)}+M_{(0)}$ (8)
    $-\left(B_{2}^{\top}Z^{(0)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(0)}C_{l,(0)}\right)^{\top}\left(R_{22}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(0)}D_{l,2}\right)^{-1}\left(B_{2}^{\top}Z^{(0)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(0)}C_{l,(0)}\right),$
    where $A_{(0)}=A-\begin{bmatrix}B_{1}&B_{2}\end{bmatrix}R(0)^{-1}S(0),\,C_{l,(0)}=C_{l}-\begin{bmatrix}D_{l,1}&D_{l,2}\end{bmatrix}R(0)^{-1}S(0),1\leq l\leq r$$M_{(0)}=Q-S^{\top}(0)R(0)^{-1}S(0)$.
  2. 2.
    For $k\geq 1$, update $P^{(k)}$ via:
@@ 2.3 Algorithm Design
    and solving $Z^{(k)}$, which is the unique stabilizing solution to the following algebraic Riccati equation with a definite quadratic term:
    $0=Z^{(k)}A_{(k)}+A_{(k)}^{\top}Z^{(k)}+\sum_{l=1}^{r}C_{l,(k)}^{\top}Z^{(k)}C_{l,(k)}+M_{(k)}$ (10)
    $-\left(B_{2}^{\top}Z^{(k)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(k)}C_{l,(k)}\right)^{\top}\left(R_{22}(P^{(k)})+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(k)}D_{l,2}\right)^{-1}\left(B_{2}^{\top}Z^{(k)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(k)}C_{l,(k)}\right).$
@@
    where
    $R^{\sharp}_{22}(P)=R_{11}(P)-R_{12}(P)R_{22}(P)^{-1}R_{21}(P),\begin{bmatrix}N_{1,(k-1)}\\ N_{2,(k-1)}\end{bmatrix}=\begin{bmatrix}B_{1}^{\top}Z^{(k-1)}+\sum_{l=1}^{r}D_{l,1}^{\top}Z^{(k-1)}C_{l,(k-1)}\\ B_{2}^{\top}Z^{(k-1)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(k-1)}C_{l,(k-1)}\end{bmatrix}$. (12)
Under some natural assumptions, we will show:
  • The sequences $\{Z^{(k)}\}$ and $\{P^{(k)}\}$ are well-defined. Each algebraic Riccati equation admits a unique stabilizing solution $\tilde{P}_{L}\succeq Z^{(k)}\succeq 0(k=0,1,2,3\cdots)$, where $\tilde{P}_{L}$ be the stabilizing solution of the ARE (7).
  • These sequences are convergent and we have
[…]
@@ 3 The Main Results
3.1 Analysis of Structural Characteristics for ZSLQSDG
From stochastic GTAREs (3), we can define a mapping $\mathcal{G}:\mathrm{Dom}\,\mathcal{G}\to\mathbb{S}^{n}$ that satisfies the following relation:
@@ 3.1 Analysis of Structural Characteristics for ZSLQSDG
The nonlinear function $\mathcal{G}$ is not defined over the entire space $\mathbb{S}^{n}$. Instead, it is well-defined on the subset:
@@
The aforementioned mapping $\mathcal{G}$ serves as the fundamental foundation for constructing the algorithm. We begin by detailing its general properties, which collectively ensure the algorithm’s ability to operate iteratively.
@@
$\begin{bmatrix}K_{1}(P+Z)\\ K_{2}(P+Z)\end{bmatrix}=\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}-R(P+Z)^{-1}\begin{bmatrix}B_{1}^{\top}Z+\sum_{l=1}^{r}D_{l,1}^{\top}Z\bigl(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P)\bigr)\\ B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z\bigl(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P)\bigr)\end{bmatrix}$,
where
$\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}=-R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}$.
@@
Define the increment,
$\Delta=\begin{bmatrix}K_{1}(P+Z)\\ K_{2}(P+Z)\end{bmatrix}-\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}$.
By leveraging the definitions of $K_{1}(\cdot),K_{2}(\cdot)$, and multiplying both sides of the equation by $R(P+Z)$, we obtain:
@@
Thus,
$R(P+Z)\Delta=\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}+\Delta R\cdot R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}-\begin{bmatrix}S_{1}(P+Z)\\ S_{2}(P+Z)\end{bmatrix}$.
Using the identity,
$R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}=-\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}$,
we obtain:
$R(P+Z)\Delta=-\begin{bmatrix}B_{1}^{\top}Z+\sum_{l=1}^{r}D_{l,1}^{\top}Z\bigl(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P)\bigr)\\ B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z\bigl(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P)\bigr)\end{bmatrix}$.
Multiplying both sides by $R(P+Z)^{-1}$ yields:
$\begin{bmatrix}K_{1}(P+Z)\\ K_{2}(P+Z)\end{bmatrix}-\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}=-R(P+Z)^{-1}\begin{bmatrix}B_{1}^{\top}Z+\sum_{l=1}^{r}D_{l,1}^{\top}Z\bigl(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P)\bigr)\\ B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z\bigl(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P)\bigr)\end{bmatrix}$.
Proposition 3.2 ([*]Dragan2013book).
@@
$\mathcal{G}(P)={}$ $P(A+B_{1}\varTheta_{1}+B_{2}\varTheta_{2})+(A+B_{1}\varTheta_{1}+B_{2}\varTheta_{2})^{\top}P$
$+\sum_{l=1}^{r}(C_{l}+D_{l,1}\varTheta_{1}+D_{l,2}\varTheta_{2})^{\top}P(C_{l}+D_{l,1}\varTheta_{1}+D_{l,2}\varTheta_{2})+Q$
$+\begin{bmatrix}\varTheta_{1}\\ \varTheta_{2}\end{bmatrix}^{\top}R\begin{bmatrix}\varTheta_{1}\\ \varTheta_{2}\end{bmatrix}+\begin{bmatrix}\varTheta_{1}\\ \varTheta_{2}\end{bmatrix}^{\top}\begin{bmatrix}S_{1}\\ S_{2}\end{bmatrix}+\begin{bmatrix}S_{1}\\ S_{2}\end{bmatrix}^{\top}\begin{bmatrix}\varTheta_{1}\\ \varTheta_{2}\end{bmatrix}$
$-\begin{bmatrix}K_{1}(P)-\varTheta_{1}\\ K_{2}(P)-\varTheta_{2}\end{bmatrix}^{\top}R(P)\begin{bmatrix}K_{1}(P)-\varTheta_{1}\\ K_{2}(P)-\varTheta_{2}\end{bmatrix},\,\text{where}\,\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}=-R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}$.
@@
Let $P$, $Z$ satisfy $P,P+Z\in\mathrm{Dom}\,\mathcal{G}$. Denote the feedback gain matrix as:
$K(P)\triangleq\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}=-R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}$,
and the auxiliary matrix as:
$N(P,Z)\triangleq\begin{bmatrix}N_{1}(P,Z)\\ N_{2}(P,Z)\end{bmatrix}=\begin{bmatrix}B_{1}^{\top}Z+\sum_{l=1}^{r}D_{l,1}^{\top}Z(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P))\\ B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P))\end{bmatrix}$. (14)
Then the following identity holds:
@@
$+\sum_{l=1}^{r}(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P))^{\top}Z(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P))-N(P,Z)^{\top}R(P+Z)^{-1}N(P,Z).$
@@
Starting from the definition of $\mathcal{G}$,
$\mathcal{G}(P+Z)=(P+Z)A+A^{\top}(P+Z)+\sum_{l=1}^{r}C_{l}^{\top}(P+Z)C_{l}+Q-S(P+Z)^{\top}R(P+Z)^{-1}S(P+Z).$
From Proposition 3.2, expanding and rearranging terms, we obtain:
$\mathcal{G}(P+Z)=$ $PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+Q+K(P)^{\top}R(P)K(P)+S(P)^{\top}K(P)+K(P)^{\top}S(P)$
@@
Thus,
$-N(P,Z)^{\top}R(P+Z)^{-1}N(P,Z).$
3.2 A Criteria for the Existence and Uniqueness of Stabilizing Solutions of AREs with Definite Quadratic Term
Based on the structure of $\mathcal{G}$ and Proposition 3.3, a series of interconnected sub-algebraic Riccati equations with definite quadratic terms can be constructed, as described in Section 2.3. A key requirement for the algorithm is the existence and uniqueness of stabilizing solutions to these equations. We present a criterion for this purpose in what follows.
Definition 3.4 ([*]Dragan2013book).
@@
$\begin{cases}dX(t)=AX(t)dt+\sum_{l=1}^{r}C_{l}X(t)dw_{l}(t)\\ dY(t)=E_{0}X(t)dt+\sum_{l=1}^{r}E_{l}X(t)dw_{l}(t)\end{cases},$ (15)
where $E_{l}\in R^{q\times n}$ for $0\leq l\leq r$, we denote this system by $\left[E_{0},E_{1},\dots,E_{r};A,C_{1},\dots,C_{r}\right]$. It is said to be stochastically detectable if there exists a constant matrix $\varTheta\in\mathbb{R}^{n\times q}$ such that the system
$\left(A+\varTheta E_{0},\ C_{1}+\varTheta E_{1},\ \dots,\ C_{r}+\varTheta E_{r}\right)$
is mean-square stable.
@@
If the system $\left[E_{0},E_{1},\dots,E_{r};A,C_{1},\dots,C_{r}\right]$ is stochastically detectable, then the following statements are equivalent:
  • (a.)
    The system $(A,C_{1},\dots,C_{r})$ is mean-square stable. Let $\mathcal{L}^{*}$ denote the linear operator associated with this system, defined by
    $\mathcal{L}^{*}(P)=PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l},\quad\forall P\in\mathbb{S}^{n}.$
    Then all eigenvalues of $\mathcal{L}^{*}$ lie in the left half-plane, i.e., $\operatorname{Spec}\mathcal{L}^{*}\subset\mathbb{C}^{-}$.
  • (b.)
    There exists a matrix $P\in\overline{\mathbb{S}}_{+}^{n}$ satisfying the matrix equation:
    $PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+\sum_{l=0}^{r}E_{l}^{\top}E_{l}=0.$
@@
Proof of (a.) $\Rightarrow$ (b.): Since $\sum_{l=0}^{r}E_{l}^{\top}E_{l}\succeq 0$ and the system $(A,C_{1},\dots,C_{r})$ is mean-square stable, it follows from Theorem 3.2.2. and Theorem 2.7.5 in [12] that the equation
$PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+\sum_{l=0}^{r}E_{l}^{\top}E_{l}=0$
has a solution $P\in\overline{\mathbb{S}}_{+}^{n}$.
Proof of (b.) $\Rightarrow$ (a.): This direction follows directly from Remark 4.1.5 in [12]. For a detailed argument, see Theorem 4.1.7 and Remark 4.1.5 in the same reference. Furthermore, if the system $(A,C_{1},\dots,C_{r})$ is mean-square stable, then $\operatorname{Spec}\mathcal{L}^{*}\subset\mathbb{C}^{-}$; this is a consequence of Theorem 3.2.2. and Theorem 2.7.7 in [12]. ∎
Now we consider AREs of the form:
$PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+Q-\left(B^{\top}P+\sum_{l=1}^{r}D_{l}^{\top}PC_{l}\right)^{\top}\left(R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l}\right)^{-1}\left(B^{\top}P+\sum_{l=1}^{r}D_{l}^{\top}PC_{l}\right)=0,$ (16)
@@
Define the operator $\bm{\Lambda}^{[A,C_{1},\dots,C_{r};B,D_{1},\dots,D_{r};Q,R]}:\mathbb{S}^{n}\rightarrow\mathbb{S}^{n+m}$ associated with the ARE (16) as:
$\bm{\Lambda}^{[A,C_{1},\dots,C_{r};B,D_{1},\dots,D_{r};Q,R]}(P)=\begin{bmatrix}PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+Q&\Big(B^{\top}P+\sum_{l=1}^{r}D_{l}^{\top}PC_{l}\Big)^{\top}\\ B^{\top}P+\sum_{l=1}^{r}D_{l}^{\top}PC_{l}&R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l}\end{bmatrix}$. (17)
Define the set $\bm{\Gamma}^{[A,C_{1},\dots,C_{r};B,D_{1},\dots,D_{r};Q,R]}$ related to ARE (16) as:
$\bm{\Gamma}^{[A,C_{1},\dots,C_{r};B,D_{1},\dots,D_{r};Q,R]}=\left\{P\in\mathbb{S}^{n}\,\middle|\,\bm{\Lambda}(P)\succeq 0,\ R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l}\succ 0\right\}.$ (18)
Definition 3.6 ([*]Dragan2013book).
The system (16) is denoted by $\left[A,C_{1},\dots,C_{r};B,D_{1},\dots,D_{r}\right]$ is said to be stochastically stabilizable if there exists a constant matrix $\varTheta\in\mathbb{R}^{n\times m}$ such that the system
$\left(A+B\varTheta,C_{1}+D_{1}\varTheta,\dots,C_{r}+D_{r}\varTheta\right)$
is mean-square stable.

### v2

@@ 2.2 Zero-Sum Linear-Quadratic Stochastic Differential Games in Infinite Horizons
2.3 Nested Iterative Scheme
In this paper, we aim to solve for the closed-loop saddle points numerically for the $({SDG})^{\,0}_{\infty}$ problem with given parameter. Consider the sequences $\{P^{(k)}\}_{k\geq 0}$ and $\{Z^{(k)}\}_{k\geq 0}$ constructed according to the following procedure:
1. For $k=0$, set $P^{(0)}=0$, and let $Z^{(0)}$ be the unique stabilizing solution to the following ARE with sign-definite quadratic term:
$Z^{(0)}A_{(0)}+A_{(0)}^{\top}Z^{(0)}+\sum_{l=1}^{r}C_{l,(0)}^{\top}Z^{(0)}C_{l,(0)}+M_{(0)}-(B_{2}^{\top}Z^{(0)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(0)}C_{l,(0)})^{\top}$ (6)
$\quad\times(R_{22}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(0)}D_{l,2})^{-1}(B_{2}^{\top}Z^{(0)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(0)}C_{l,(0)})=0,$
where $A_{(0)}=A-\begin{bmatrix}B_{1}&B_{2}\end{bmatrix}R(0)^{-1}S(0),\,\,C_{l,(0)}=C_{l}-\begin{bmatrix}D_{l,1}&D_{l,2}\end{bmatrix}R(0)^{-1}S(0)(1\leq l\leq r)$,   $M_{(0)}=Q-S^{\top}(0)R(0)^{-1}S(0)$.
2. For $k\geq 1$, update $P^{(k)}$ via:
@@
and solving $Z^{(k)}$, which is the unique stabilizing solution to the following ARE with sign-definite quadratic term:
$Z^{(k)}A_{(k)}+A_{(k)}^{\top}Z^{(k)}+\sum_{l=1}^{r}C_{l,(k)}^{\top}Z^{(k)}C_{l,(k)}+M_{(k)}-(B_{2}^{\top}Z^{(k)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(k)}C_{l,(k)})^{\top}$ (8)
$\quad\times(R_{22}(P^{(k)})+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(k)}D_{l,2})^{-1}(B_{2}^{\top}Z^{(k)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(k)}C_{l,(k)})=0,$
@@
where $R^{\sharp}_{22}(P)=R_{11}(P)-R_{12}(P)R_{22}(P)^{-1}R_{21}(P),$ (10)
and $\begin{bmatrix}N_{1,(k-1)}\\ N_{2,(k-1)}\end{bmatrix}=\begin{bmatrix}B_{1}^{\top}Z^{(k-1)}+\sum_{l=1}^{r}D_{l,1}^{\top}Z^{(k-1)}C_{l,(k-1)}\\ B_{2}^{\top}Z^{(k-1)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(k-1)}C_{l,(k-1)}\end{bmatrix}$.
In the subsequent content, we will show:
  • The sequences $\{P^{(k)}\}_{k\geq 0}$ and $\{Z^{(k)}\}_{k\geq 0}$ are well-defined. The AREs in (6) and (8) admits unique stabilizing solutions $Z^{(k)}\succeq 0(k=0,1,2,\dotsb)$ such that $\tilde{P}_{L}\succeq\sum^{h}_{k=0}Z^{(k)}(h=0,1,2,\dotsb)$, where $\tilde{P}_{L}$ be the stabilizing solution of the corresponding ARE (5).
  • These sequences are convergent and we have $\lim_{k\to\infty}P^{(k)}=\tilde{P}$, where $\tilde{P}$ is the stabilizing solution to the stochastic GTARE (3).
@@ 3 The Main Results
3.1 Analysis of Structural Characteristics for $({SDG})^{\,0}_{\infty}$
From stochastic GTARE (3), we can define a mapping $\mathcal{G}:\mathrm{Dom}\,\mathcal{G}\to\mathbb{S}^{n}$ that satisfies the following relation:
@@ 3.1 Analysis of Structural Characteristics for $({SDG})^{\,0}_{\infty}$
The nonlinear function $\mathcal{G}$ is not defined over the entire space $\mathbb{S}^{n}$. Instead, it is well-defined on the subset
@@
The aforementioned mapping $\mathcal{G}$ serves as the fundamental foundation for constructing the nested iterative scheme. We begin by detailing its general properties, which collectively ensure the computational method’s ability to operate iteratively.
First, we define two operators that will be used in the subsequent content:
$\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}=-R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix},\begin{bmatrix}N_{1}(P,Z)\\ N_{2}(P,Z)\end{bmatrix}=\begin{bmatrix}B_{1}^{\top}Z+\sum_{l=1}^{r}D_{l,1}^{\top}Z\bigl(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P)\bigr)\\ B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z\bigl(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P)\bigr)\end{bmatrix}$, (12)
where $B_{i},D_{l,i}(i=1,2;1\leq l\leq r)$ is defined in (1), and $R(P),S_{1}(P),S_{2}(P)$ is defined in (4).
@@
$\begin{bmatrix}K_{1}(P+Z)\\ K_{2}(P+Z)\end{bmatrix}=\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}-R(P+Z)^{-1}\begin{bmatrix}N_{1}(P,Z)\\ N_{2}(P,Z)\end{bmatrix}$.
@@
Define the increment $\Delta=\begin{bmatrix}K_{1}(P+Z)\\ K_{2}(P+Z)\end{bmatrix}-\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}$. By leveraging the definitions of $K_{1}(\cdot),K_{2}(\cdot)$ in (12), and multiplying both sides of the equation by $R(P+Z)$, we obtain:
@@
Using the identity, $R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}=-\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}$, we have
$R(P+Z)\Delta=\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}+\Delta R\cdot R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}-\begin{bmatrix}S_{1}(P+Z)\\ S_{2}(P+Z)\end{bmatrix}=-\begin{bmatrix}N_{1}(P,Z)\\ N_{2}(P,Z)\end{bmatrix}$.
Multiplying both sides by $R(P+Z)^{-1}$ yields the relation. ∎
Proposition 3.2 ([Dragan2013book]).
@@
$\mathcal{G}(P)=P(A+B_{1}\varTheta_{1}+B_{2}\varTheta_{2})+(A+B_{1}\varTheta_{1}+B_{2}\varTheta_{2})^{\top}P+Q$
$+\sum_{l=1}^{r}(C_{l}+D_{l,1}\varTheta_{1}+D_{l,2}\varTheta_{2})^{\top}P(C_{l}+D_{l,1}\varTheta_{1}+D_{l,2}\varTheta_{2})+\begin{bmatrix}\varTheta_{1}\\ \varTheta_{2}\end{bmatrix}^{\top}R(0)\begin{bmatrix}\varTheta_{1}\\ \varTheta_{2}\end{bmatrix}$
$+\begin{bmatrix}\varTheta_{1}\\ \varTheta_{2}\end{bmatrix}^{\top}\begin{bmatrix}S_{1}\\ S_{2}\end{bmatrix}+\begin{bmatrix}S_{1}\\ S_{2}\end{bmatrix}^{\top}\begin{bmatrix}\varTheta_{1}\\ \varTheta_{2}\end{bmatrix}-\begin{bmatrix}K_{1}(P)-\varTheta_{1}\\ K_{2}(P)-\varTheta_{2}\end{bmatrix}^{\top}R(P)\begin{bmatrix}K_{1}(P)-\varTheta_{1}\\ K_{2}(P)-\varTheta_{2}\end{bmatrix}$.
@@
Let $P$, $Z$ satisfy $P,P+Z\in\mathrm{Dom}\,\mathcal{G}$. Then the following identity holds:
@@
$-N(P,Z)^{\top}R(P+Z)^{-1}N(P,Z),\quad\text{where}\,\,N(P,Z)=\begin{bmatrix}N_{1}(P,Z)\\ N_{2}(P,Z)\end{bmatrix}$.
@@
Starting from the definition of $\mathcal{G}$, $\mathcal{G}(P+Z)=(P+Z)A+A^{\top}(P+Z)+\sum_{l=1}^{r}C_{l}^{\top}(P+Z)C_{l}+Q-S(P+Z)^{\top}R(P+Z)^{-1}S(P+Z).$
From Proposition 3.2, expanding and rearranging terms, we obtain $\mathcal{G}(P+Z)=$
$PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+Q+K(P)^{\top}R(P)K(P)+S(P)^{\top}K(P)+K(P)^{\top}S(P)$
@@
Substitute the above term into the rearranged expression of $\mathcal{G}(P+Z)$, and the target identity is thus verified to hold. ∎
3.2 Existence and Uniqueness of Stabilizing Solution for ARE with Sign-Definite Quadratic Term
Based on the structure of $\mathcal{G}$ and Proposition 3.3, a series of interconnected sub-algebraic Riccati equations with sign-definite quadratic terms can be constructed, as detailed in Section 2.3. In each iteration step, the existence and uniqueness of stabilizing solutions to these equations are of particular importance. In what follows, we establish a criterion for such equations and separately present some related properties—both of which will be repeatedly employed in proving the iteration sequence’s convergence.
Definition 3.4 ([Dragan2013book]).
@@ 3.2 Existence and Uniqueness of Stabilizing Solution for ARE with Sign-Definite Quadratic Term
$\begin{cases}dX(t)=AX(t)dt+\sum_{l=1}^{r}C_{l}X(t)dw_{l}(t)\\ dY(t)=E_{0}X(t)dt+\sum_{l=1}^{r}E_{l}X(t)dw_{l}(t)\end{cases},$
where $E_{l}\in R^{q\times n}(0\leq l\leq r)$, we denote this system by $\left[E_{0},E_{1},\dotsb,E_{r};A,C_{1},\dotsb,C_{r}\right]$. It is said to be stochastically detectable if there exists a constant matrix $\varTheta\in\mathbb{R}^{n\times q}$ such that the system $(A+\varTheta E_{0},\ C_{1}+\varTheta E_{1},\dotsb,C_{r}+\varTheta E_{r})$ is mean-square stable.
@@
If the system $\left[E_{0},E_{1},\dotsb,E_{r};A,C_{1},\dotsb,C_{r}\right]$ is stochastically detectable, then the following statements are equivalent:
  • (a).
    The system $(A,C_{1},\dotsb,C_{r})$ is mean-square stable. Let $\mathcal{L}^{*}$ denote the linear operator associated with this system, defined by $\mathcal{L}^{*}(P)=PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l};\,\forall P\in\mathbb{S}^{n}$. Then all eigenvalues of $\mathcal{L}^{*}$ lie in the left half-plane, i.e., $\operatorname{Spec}\mathcal{L}^{*}\subset\mathbb{C}^{-}$.
  • (b).
    There exists a matrix $P\in\overline{\mathbb{S}}_{+}^{n}$ satisfying the matrix equation $PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+\sum_{l=0}^{r}E_{l}^{\top}E_{l}=0$.
@@
(a) $\Rightarrow$ (b): Since $\sum_{l=0}^{r}E_{l}^{\top}E_{l}\succeq 0$ and the system $(A,C_{1},\dotsb,C_{r})$ is mean-square stable, it follows from Theorem 3.2.2. and Theorem 2.7.5 in [Dragan2013book] that the equation $PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+\sum_{l=0}^{r}E_{l}^{\top}E_{l}=0$ has a solution $P\in\overline{\mathbb{S}}_{+}^{n}$.
(b) $\Rightarrow$ (a): This conclusion follows directly from Remark 4.1.5 in [Dragan2013book]. For a detailed argument, see Theorem 4.1.7 and Remark 4.1.5 in the same reference. Furthermore, if the system $(A,C_{1},\dotsb,C_{r})$ is mean-square stable, then $\operatorname{Spec}\mathcal{L}^{*}\subset\mathbb{C}^{-}$. This is a consequence of Theorem 3.2.2. and Theorem 2.7.7 in [Dragan2013book]. ∎
Now we consider ARE of the form
$PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+Q-(B^{\top}P+\sum_{l=1}^{r}D_{l}^{\top}PC_{l})^{\top}(R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l})^{-1}(B^{\top}P+\sum_{l=1}^{r}D_{l}^{\top}PC_{l})=0,$ (13)
@@
Define the operator $\bm{\Lambda}^{[A,C_{1},\dotsb,C_{r};B,D_{1},\dotsb,D_{r};Q,R]}:\mathbb{S}^{n}\rightarrow\mathbb{S}^{n+m}$ associated with the ARE (13) as
$\bm{\Lambda}^{[A,C_{1},\dotsb,C_{r};B,D_{1},\dotsb,D_{r};Q,R]}(P)=\begin{bmatrix}PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+Q&PB+\sum_{l=1}^{r}C_{l}^{\top}PD_{l}\\ B^{\top}P+\sum_{l=1}^{r}D_{l}^{\top}PC_{l}&R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l}\end{bmatrix}$.
Define the set $\bm{\Gamma}^{[A,C_{1},\dotsb,C_{r};B,D_{1},\dotsb,D_{r};Q,R]}$ related to ARE (13) as
$\left\{P\in\mathbb{S}^{n}\,\middle|\,\bm{\Lambda}^{[A,C_{1},\dotsb,C_{r};B,D_{1},\dotsb,D_{r};Q,R]}(P)\succeq 0,\ R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l}\succ 0\right\}.$
Definition 3.6 ([Dragan2013book]).
The system associated with the ARE (13) denoted as $A,C_{1},\dotsb,C_{r};\,B,D_{1},\\ \dotsb,D_{r}]$ is said to be mean-square stabilizable if there exists a constant matrix $\varTheta\in\mathbb{R}^{n\times m}$ such that the system $\left(A+B\varTheta,C_{1}+D_{1}\varTheta,\dotsb,C_{r}+D_{r}\varTheta\right)$ is mean-square stable.